# An algorithmic version of the Hajnal–Szemerédi theorem

### v2

@@ 6.2. Ideal case
Proof of Lemma 6.3..
Note that a $K_{r-s}$-factor of $G[B^{\prime}]$ has size $\frac{|B^{\prime}|}{r-s}$. We arbitrarily fix a $K_{r-s}$-factor $\mathcal{T}$ of $G[B^{\prime}]$. We contract all members of $\mathcal{T}$ individually to a set of disjoint vertices, denoted by $U$. For each $u\in U$, let $K_{r-s}^{u}$ be the copy of $K_{r-s}$ corresponding to $u$. Next we construct an auxiliary bipartite graph $H_{s}$ with two parts $U$ and $A_{s}^{\prime}$. For any $u\in U$ and $v\in A_{s}^{\prime}$, $uv\in E(H_{s})$ if and only if $v$ is common neighbor of vertices in $K_{r-s}^{u}$. Hence $d_{H_{s}}(u)\geq|A_{s}^{\prime}|-(r-s)\mu^{1/5}n$ and $d_{H_{s}}(v)\geq\frac{|B^{\prime}|}{r-s}-\mu^{1/5}n$, as every vertex in $K_{r-s}^{u}$ has at most $\mu^{1/5}n$ non-neighbors in $A_{s}^{\prime}$ and $v$ has at most $\mu^{1/5}n$ non-neighbors in $B^{\prime}$. Thus we can find a matching of size $|A_{s}^{\prime}|$ in time $O(n^{4})$ in $H_{s}$ by the Edmonds algorithm, which corresponds to a $K_{r-s+1}$-factor of size $|A_{s}^{\prime}|$ in $G[A_{s}^{\prime}\cup B^{\prime}]$. (Note that if $r-s=1$, then $H_{s}$ is $G[A_{s}^{\prime}\cup B^{\prime}]$.)
Iterate this procedure $s$ times we obtain a $K_{r}$-factor of size $|A_{1}^{\prime}|$ in $G[A_{1}^{\prime}\cup\dots\cup A_{s}^{\prime}\cup B^{\prime}]$. ∎
@@ 6.3. Adjusting the partition and more definitions
Note that the current partition $A_{1},\dots,A_{s}$, $B$ only captures the rough structure of $G$. Motivated by the ideal case, we need to further classify the vertices of $G$. There are two steps in this section. First we move a small number of vertices and obtain a new partition $A_{1}^{1},A_{2}^{1},\ldots,A_{s}^{1},B^{1}$ such that vertices in each $A_{i}^{1}$ have small degrees inside $A_{i}^{1}$ and have large degrees outside of $A_{i}^{1}$. However, $A_{1}^{1},A_{2}^{1},\ldots,A_{s}^{1},B^{1}$ may be unbalanced. Then by removing a $K_{r}$-tiling of small size from $G$, they will become balanced again (Lemma 6.12).
@@
[…]
@@
By Lemma 6.6 (4), every vertex $v\in B^{1}$ satisfies $d_{A_{i}}(v)\geq 2\mu^{1/3}n$ for each $i\in[s]$. However, some vertices in $B^{1}$ may have many non-neighbors in some $A_{i}$, that is, they violate the property required in Lemma 6.3. Hence we mark such vertices as bad and try to cover them by a $K_{r}$-tiling so that the remaining part of $G$ can be in the ideal case. To achieve it, we show that there are only a small number of such bad vertices (Fact 6.8).
@@
Proof of Fact 6.8.
First we consider the bad vertices in $B\cap B^{1}$. For each $i\in[s]$, note that $e(G[A_{i},B])\geq|A_{i}|(\delta(G)-(s-1)n/r)-2e(G[A_{i}])\geq|A_{i}|((r-1)n/r-c-(s-1)n/r)-2\mu n^{2}$ and $|B|=(r-s)n/r$. So we have $e(G[A_{i},B])\geq|A_{i}||B|-2\mu n^{2}-|A_{i}|c$, and thus there are at most $2\mu n^{2}+|A_{i}|c$ non-edges between $A_{i}$ and $B$. For a bad vertex $v$ in $B$, since $d_{A_{i}}(v)<|A_{i}|-2\mu^{2/5}n$, it contributes at least $2\mu^{2/5}n$ non-edges between $A_{i}$ and $B$. So the number of bad vertices in $B$ is at most $s\cdot\frac{2\mu n^{2}+|A_{i}|c}{2\mu^{2/5}n}<s\mu^{3/5}n+\frac{s\mu^{3/5}n}{2r}$ as $c\ll\mu n$. Moreover, as $|B^{1}\setminus B|<s\mu^{2/3}n$ by Fact 6.5 (1), there are at most $s\mu^{2/3}n$ bad vertices in $B^{1}\setminus B$. Therefore, there are at most $s\mu^{3/5}n+\frac{s\mu^{3/5}n}{2r}+s\mu^{2/3}n\leq s(2\mu^{3/5}+\mu^{2/3})n$ bad vertices in $B^{1}$. ∎
Let $B^{1}_{b}$ be the set of bad vertices in $B^{1}$ and $B^{1}_{g}:=B^{1}\setminus B^{1}_{b}$. Then $|B^{1}_{b}|\leq 3s\mu^{3/5}n$ and $|B^{1}_{g}|\geq(r-s)\frac{n}{r}+t_{s+1}-3s\mu^{3/5}n\geq(r-s)\frac{n}{r}-4s\mu^{3/5}n$.
@@
Thus, we have $\delta(G[B^{1}_{g}])\geq(1-\frac{1}{r-s}-5rs\mu^{3/5})|B^{1}_{g}|$. Moreover, $B^{1}_{g}$ has no $\frac{\gamma_{s+1}}{4}$-sparse $\frac{|B^{1}_{g}|}{r-s}$-set, as $\mu\ll\gamma_{s+1}$. Then we apply Fact 4.13 on $G[B^{1}_{g}]$ with $k=r-s+1$, $\varepsilon=5rs\mu^{3/5}$ and $\gamma_{H}=\frac{\gamma_{s+1}}{4}$. Since $\mu\ll\gamma_{s+1}$ and $c=o(n^{1/r^{r+3}})$, there are at least $\mu^{7/12}n^{r-s+1}$ copies of $K_{r-s+1}$ in $G[B^{1}_{g}]$. ∎
@@ 6.4. Admissibility and the characterization lemma
The following definition is key to our proofs.
@@
    A $K_{r}$-tiling $F$ is called $I$-admissible if $\textbf{s}(F)|_{i}=t_{i}$ for each $i\in I$ and $|F|\leq(3t_{I})^{r}$.
@@
    A $K_{r}$-tiling $F$ is called strong admissible if $\textbf{s}(F)|_{i}=t_{i}$ for each $i\in[s+1]$, $|F|\leq(3t_{[s+1]})^{r}+2$ and $F$ contains a copy of $K_{r}$ which contains an odd number of vertices of $B^{1}$.
The last thing we need to introduce is a set $I$ of indices for small slacks. To distinguish “large” and “small”, we use Algorithm 3. It divides $r^{\prime}$ real numbers into two groups, those with large absolute values and those with small absolute values, in a recursive manner. Assume that the algorithm outputs $I=[i_{1}^{\prime},i_{2}^{\prime}]$. Then for each $i\in I$, we have $|t_{i}|\leq(3rc/0.99)^{r^{r+1}}=(rc/0.33)^{r^{r+1}}$. For each $j\in[r^{\prime}]\setminus I$, $|t_{j}|\geq\max\{(3rt_{i_{1}^{\prime}})^{r},|3rt_{i_{2}^{\prime}}|^{r}\}$ and so $|t_{j}|\geq\max\{(3t_{I})^{r},c/0.99\}$, where $t_{I}=\sum_{i\in I}|t_{i}|$. Moreover, the running time of Algorithm 3 is $O(r^{\prime})$.
@@
5      if $t_{j}\leq\max\{(3rt_{j+1})^{r},|3rt_{\ell-1}|^{r}\}$, or $|t_{\ell}|\leq\max\{(3rt_{j+1})^{r},|3rt_{\ell-1}|^{r}\}$ then
6            Case 1: if $t_{j}\leq\max\{(3rt_{j+1})^{r},|3rt_{\ell-1}|^{r}\}$ and $|t_{\ell}|\leq\max\{(3rt_{j+1})^{r},|3rt_{\ell-1}|^{r}\}$,
7            then $I:=I\cup\{j,\ell\}$, $j:=j-1$ and $\ell:=\ell+1$;
8             Case 2: if $t_{j}\leq\max\{(3rt_{j+1})^{r},|3rt_{\ell-1}|^{r}\}$ but $|t_{\ell}|>\max\{(3rt_{j+1})^{r},|3rt_{\ell-1}|^{r}\}$,
9            then $I:=I\cup\{j\}$ and $j:=j-1$;
10             Case 3: if $|t_{\ell}|\leq\max\{(3rt_{j+1})^{r},|3rt_{\ell-1}|^{r}\}$ but $t_{j}>\max\{(3rt_{j+1})^{r},|3rt_{\ell-1}|^{r}\}$,
@@
Now we first sort $t_{1},t_{2},\ldots,t_{s+1},\frac{c}{0.99},-\frac{c}{0.99}$ by size and obtain that $t_{\tau(1)}\geq\cdots\geq t_{\tau(i_{1})}>\frac{c}{0.99}\geq t_{\tau(i_{1}+1)}\geq\cdots\geq t_{\tau(i_{2})}\geq-\frac{c}{0.99}>t_{\tau(i_{2}+1)}\geq\cdots\geq t_{\tau(s+1)}$ in time $O(r^{2})$. Then we apply Algorithm 3 with $r^{\prime}=s+1$ on $t_{\tau(1)},\ldots,t_{\tau(s+1)}$ and output $I=[i_{1}^{\prime},i_{2}^{\prime}]$ in time $O(s)$, with the following properties. Let $t_{I}:=\sum_{i\in I}|t_{\tau(i)}|$ and $I^{*}:=\{\tau(i):i\in I\}$. We have
@@
    for $j\in[r^{\prime}]\setminus I$, $|t_{\tau(j)}|\geq\max\{(3rt_{\tau(i_{1}^{\prime})})^{r},|3rt_{\tau(i_{2}^{\prime})}|^{r}\}$ and so $|t_{\tau(j)}|\geq\max\{(3t_{I})^{r},c/0.99\}$,
@@
Suppose $\mu>0$ and $1/n\ll\mu$. Given an $n$-vertex graph $G$ with $\delta(G)\geq(1-1/r)n-c$, let $(A_{1}^{1},\dots,A_{s}^{1},B^{1})$ be the partition defined above in this section and suppose $s<r$. Let $I,I^{*}\subseteq[s+1]$ and $B_{b}^{1}\subseteq B^{1}$ be defined as above. Then the following holds.
@@
    If $r-s\neq 2$, $I\neq[s+1]$, $B_{b}^{1}\neq\emptyset$, or $B^{1}$ is not disconnected with two odd components, then $G$ has a $K_{r}$-factor if and only if $G$ contains an $I^{*}$-admissible $K_{r}$-tiling.
@@
Moreover, the existence of such $K_{r}$-tiling in $G$ can be checked by an algorithm that runs in time $n^{O(1)}2^{O(c^{{r}^{r+2}})}$, and if $G$ has a $K_{r}$-factor, it also outputs a $K_{r}$-factor.

### v3

@@ 6.2. Ideal case
Proof.
Note that a $K_{r-s}$-factor of $G[B^{\prime}]$ has size $\frac{|B^{\prime}|}{r-s}$. We arbitrarily fix a $K_{r-s}$-factor $\mathcal{T}$ of $G[B^{\prime}]$. We contract all members of $\mathcal{T}$ individually to a set of disjoint vertices, denoted by $U$. For each $u\in U$, let $K_{r-s}^{u}$ be the copy of $K_{r-s}$ corresponding to $u$. Next we construct an auxiliary bipartite graph $H_{s}$ with two parts $U$ and $A_{s}^{\prime}$. Note that $|U|=|A_{s}^{\prime}|=\frac{|B^{\prime}|}{r-s}$. For any $u\in U$ and $v\in A_{s}^{\prime}$, $uv\in E(H_{s})$ if and only if $v$ is a common neighbor of vertices in $K_{r-s}^{u}$. Hence $d_{H_{s}}(u)\geq|A_{s}^{\prime}|-(r-s)\mu^{1/5}n>|A^{\prime}_{s}|/2$ and $d_{H_{s}}(v)\geq\frac{|B^{\prime}|}{r-s}-\mu^{1/5}n>|A^{\prime}_{s}|/2$, as every vertex in $K_{r-s}^{u}$ has at most $\mu^{1/5}n$ non-neighbors in $A_{s}^{\prime}$ and $v$ has at most $\mu^{1/5}n$ non-neighbors in $B^{\prime}$. Thus $\delta(H_{s})>|A^{\prime}_{s}|/2$ and so there is a perfect matching in $H_{s}$, i.e. a matching of size $|A_{s}^{\prime}|$. We can find the matching in time $O(n^{4})$ by the Edmonds algorithm, which corresponds to a $K_{r-s+1}$-factor of size $|A_{s}^{\prime}|$ in $G[A_{s}^{\prime}\cup B^{\prime}]$. (Note that if rs=1r-s=1, then $H_{s}$ is $G[A_{s}^{\prime}\cup B^{\prime}]$.)
Iterating this procedure $s$ times we obtain a $K_{r}$-factor of size $|A_{1}^{\prime}|$ in $G[A_{1}^{\prime}\cup\dots\cup A_{s}^{\prime}\cup B^{\prime}]$. ∎
@@ 6.3. Adjusting the partition and more definitions
Note that the current partition $A_{1},\dots,A_{s}$, $B$ only captures the rough structure of $G$. Motivated by the ideal case, we need to further classify the vertices of $G$. In this section, we move a small number of vertices and obtain a new partition $A_{1}^{1},A_{2}^{1},\ldots,A_{s}^{1},B^{1}$ such that vertices in each $A_{i}^{1}$ have small degrees inside $A_{i}^{1}$ and have large degrees outside $A_{i}^{1}$. However, $A_{1}^{1},A_{2}^{1},\ldots,A_{s}^{1},B^{1}$ may be unbalanced. Then by removing a $K_{r}$-tiling of small size from $G$, they will become balanced again (Lemmas 6.12 and 6.13).
@@
Now we define the vertex partition $A_{i}^{1}:=\{v:v~\text{is an}~A_{i}\text{-good vertex}\}$ for each $i\in[s]$ and $B^{1}:=V(G)\setminus\cup_{i=1}^{s}A_{i}^{1}$. Then $A_{1}^{1},A_{2}^{1},\ldots,A_{s}^{1},B^{1}$ may be unbalanced. We assume that $|A_{i}^{1}|=n/r+t_{i}$ for each $i\in[s]$ and $|B^{1}|=(r-s)n/r+t_{s+1}$. Note that $\sum_{i=1}^{s+1}t_{i}=0$. By Definition 6.4 and Fact 6.5, we have the following result.
@@
Proof.
(1) Since $\delta(G)\geq(1-1/r)n-c$, for any $v\in V(G)$, it has at most $\frac{n}{r}+c$ non-neighbors. Thus $\delta(G[A_{i}^{1}])\geq|A_{i}^{1}|-(\frac{n}{r}+c)\geq t_{i}-c$ and $\delta(G[B^{1}])\geq|B^{1}|-(\frac{n}{r}+c)\geq(r-s-1)n/r+t_{s+1}-c$.
(2) By Fact 6.5 (1) and (2), for each $i\in[s]$, we have $\mu^{2/3}n<t_{i}<4\mu rn+2c$. Then
$4s\mu rn-2sc<t_{s+1}=-\sum_{i=1}^{s}t_{i}<s\mu^{2/3}n.$
(3) By Definition 6.4 and Fact 6.5 (2), we have
$\Delta(G[A_{i}^{1}])<2\mu^{1/3}n+4\mu rn+2c.$
(4) By Definition 6.4 and the definition of $B^{1}$, any vertex $v\in B^{1}$ satisfies $d_{A_{i}}(v)\geq 2\mu^{1/3}n$ for each $i\in[s]$.
(5) By Lemma 6.1, we know that $B$ has no $\lambda$-sparse $\frac{n}{r}$-set of $G$, that is, for any $W\subseteq B$ with $|W|=\frac{n}{r}$, $e(G[W])\geq\lambda n^{2}$. Then for any $\frac{n}{r}$-set $U\subseteq B^{1}$, we have
$e(G[U])\geq\lambda n^{2}-2|U|\cdot(s\mu^{2/3}n)\geq\lambda n^{2}-2s\mu^{2/3}n^{2}/r\geq\frac{\lambda}{2}n^{2},$
as $\mu\ll\lambda$. Thus $B^{1}$ has no $\frac{\lambda}{2}$-sparse $\frac{n}{r}$-set of $G$. ∎
By Lemma 6.6 (4), every vertex $v\in B^{1}$ satisfies $d_{A_{i}}(v)\geq 2\mu^{1/3}n$ for each $i\in[s]$. However, some vertices in $B^{1}$ may have many non-neighbors in some $A_{i}$, that is, they violate the property required in Lemma 6.3. Hence we mark such vertices as bad and try to cover them by a $K_{r}$-tiling (Lemma 6.14) so that the remaining part of $G$ can be in the ideal case. Lemma 6.14 can cover only a small number of such bad vertices. Fact 6.8 shows that the number of such bad vertices is indeed small.
@@
Proof.
First we consider the bad vertices in $B\cap B^{1}$. For each $i\in[s]$, note that $e(G[A_{i},B])\geq|A_{i}|(\delta(G)-(s-1)n/r)-2e(G[A_{i}])\geq|A_{i}|((r-1)n/r-c-(s-1)n/r)-2\mu n^{2}$ and |B|=(rs)n/r|B|=(r-s)n/r. So we have $e(G[A_{i},B])\geq|A_{i}||B|-2\mu n^{2}-|A_{i}|c$, and thus there are at most $2\mu n^{2}+|A_{i}|c$ non-edges between $A_{i}$ and $B$. For a bad vertex $v$ in $B$, since $d_{A_{i}}(v)<|A_{i}|-2\mu^{2/5}n$, it contributes at least $2\mu^{2/5}n$ non-edges between $A_{i}$ and $B$. So the number of bad vertices in $B$ is at most $s\cdot\frac{2\mu n^{2}+|A_{i}|c}{2\mu^{2/5}n}<s\mu^{3/5}n+\frac{s\mu^{3/5}n}{2r}$ as $c\ll\mu n$. Moreover, as $|B^{1}\setminus B|<s\mu^{2/3}n$ by Fact 6.5 (1), there are at most $s\mu^{2/3}n$ bad vertices in $B^{1}\setminus B$. Therefore, there are at most $s\mu^{3/5}n+\frac{s\mu^{3/5}n}{2r}+s\mu^{2/3}n\leq s(2\mu^{3/5}+\mu^{2/3})n$ bad vertices in $B^{1}$. ∎
Let $B^{1}_{b}$ be the set of bad vertices in $B^{1}$ and $B^{1}_{g}:=B^{1}\setminus B^{1}_{b}$. Then $|B^{1}_{b}|\leq 3s\mu^{3/5}n$ and $|B^{1}_{g}|\geq(r-s)\frac{n}{r}+t_{s+1}-3s\mu^{3/5}n\geq(r-s)\frac{n}{r}-4s\mu^{3/5}n$. The next lemma gives a property of $B^{1}_{g}$ that will be used in the proof of Lemma 6.13.
@@
Thus, we have $\delta(G[B^{1}_{g}])\geq(1-\frac{1}{r-s}-5rs\mu^{3/5})|B^{1}_{g}|$. Moreover, $B^{1}_{g}$ has no $\frac{\gamma_{s+1}}{4}$-sparse $\frac{|B^{1}_{g}|}{r-s}$-set, as $\mu\ll\gamma_{s+1}$. Then we apply Fact 4.13 on $G[B^{1}_{g}]$ with k=rs+1k=r-s+1, $\varepsilon=5rs\mu^{3/5}$ and $\gamma_{H}=\frac{\gamma_{s+1}}{4}$. Since $\mu\ll\gamma_{s+1}$ and $c=o(n^{1/r^{r+3}})$, there are at least $\mu^{7/12}n^{r-s+1}$ copies of $K_{r-s+1}$ in $G[B^{1}_{g}]$. ∎
@@ 6.4. Admissibility and the characterization lemma
The following definition is key to our proofs. Recall that $|A_{i}^{1}|=n/r+t_{i}$ for each $i\in[s]$, $|B^{1}|=(r-s)n/r+t_{s+1}$ and $\sum_{i=1}^{s+1}t_{i}=0$.
@@
    A $K_{r}$-tiling $F$ is called $I$-admissible if $\textbf{s}(F)|_{i}=t_{i}$ for each $i\in I$ and $|F|\leq g(t_{I})$.
@@
    A $K_{r}$-tiling $F$ is called strong admissible if $\textbf{s}(F)|_{i}=t_{i}$ for each $i\in[s+1]$, $|F|\leq g(t_{[s+1]})+2$ and $F$ contains a copy of $K_{r}$ which contains an odd number of vertices of $B^{1}$.
The last thing we need to introduce is a set $I^{*}$ of indices for small slacks. To distinguish “large” and “small”, we use Algorithm 3. It divides $r^{\prime}$ real numbers into two groups, those with large absolute values and those with small absolute values, in a recursive manner. Note that the running time of Algorithm 3 is $O(r^{\prime})$.
@@
5 if $t_{j}\leq\max\{g(rt_{j+1}),|g(rt_{\ell-1})|\}$, or $|t_{\ell}|\leq\max\{g(rt_{j+1}),|g(rt_{\ell-1})|\}$ then
6    Case 1: if $t_{j}\leq\max\{g(rt_{j+1}),|g(rt_{\ell-1})|\}$ and $|t_{\ell}|\leq\max\{g(rt_{j+1}),|g(rt_{\ell-1})|\}$,
7     then $I:=I\cup\{j,\ell\}$, j:=j1j:=j-1 and $\ell:=\ell+1$;
8    Case 2: if $t_{j}\leq\max\{g(rt_{j+1}),|g(rt_{\ell-1})|\}$ but $|t_{\ell}|>\max\{g(rt_{j+1}),|g(rt_{\ell-1})|\}$,
9     then $I:=I\cup\{j\}$ and j:=j1j:=j-1;
10    Case 3: if $|t_{\ell}|\leq\max\{g(rt_{j+1}),|g(rt_{\ell-1})|\}$ but $t_{j}>\max\{g(rt_{j+1}),|g(rt_{\ell-1})|\}$,
@@
Before applying Algorithm 3, we sort $t_{1},t_{2},\ldots,t_{s+1},\frac{c}{0.99},-\frac{c}{0.99}$ by size and obtain that $t_{\tau(1)}\geq\cdots\geq t_{\tau(i_{1})}>\frac{c}{0.99}\geq t_{\tau(i_{1}+1)}\geq\cdots\geq t_{\tau(i_{2})}\geq-\frac{c}{0.99}>t_{\tau(i_{2}+1)}\geq\cdots\geq t_{\tau(s+1)}$ in time $O(r^{2})$. Let $x_{i}:=t_{\tau(i)}$ for each $i\in[s+1]$. Then we apply Algorithm 3 with $r^{\prime}=s+1$ on $x_{1},\ldots,x_{s+1}$ and output $I=[i_{1}^{\prime},i_{2}^{\prime}]$ in time $O(s)$, with the following properties. Let $I^{*}:=\{\tau(i):i\in I\}$, and so $t_{I^{*}}=\sum_{i\in I^{*}}|t_{i}|$. We have
@@
    for $j\in[r^{\prime}]\setminus I^{*}$, $|t_{j}|\geq\max\{g(rt_{\tau(i_{1}^{\prime})}),|g(rt_{\tau(i_{2}^{\prime})})|\}$ and so $|t_{j}|\geq\max\{g(t_{I^{*}}),c/0.99\}$,
@@
Suppose $\mu>0$ and $1/n\ll\mu$. Given an $n$-vertex graph $G$ with $\delta(G)\geq(1-1/r)n-c$, let $(A_{1}^{1},\dots,A_{s}^{1},B^{1})$ be the partition defined in Section 6.3 and suppose $s<r$. Let $I^{*}\subseteq[s+1]$ and $B_{b}^{1}\subseteq B^{1}$ be defined as above. Then the following holds.
@@
    If $r-s\neq 2$, $I^{*}\neq[s+1]$, $B_{b}^{1}\neq\emptyset$, or $B^{1}$ is not disconnected with two odd components, then $G$ has a $K_{r}$-factor if and only if $G$ contains an $I^{*}$-admissible $K_{r}$-tiling.
@@
Moreover, the existence of such $K_{r}$-tiling in $G$ can be checked by an algorithm that runs in time $n^{O(1)}2^{O(c^{{r}^{r+2}})}$ for $r\geq 4$ (resp. $n^{O(1)}2^{O(c)}$ for $r=3$), and if $G$ has a $K_{r}$-factor, it also outputs a $K_{r}$-factor.